# VGS-ATD: Robust Distributed Learning for Multi-Label Medical Image Classification Under Heterogeneous and Imbalanced Conditions

### v1

@@ Abstract
In recent years, advanced deep learning architectures have shown strong performance in medical imaging tasks. However, the traditional centralized learning paradigm poses serious privacy risks as all data is collected and trained on a single server. To mitigate this challenge, decentralized approaches such as federated learning and swarm learning have emerged, allowing model training on local nodes while sharing only model weights. While these methods enhance privacy, they struggle with heterogeneous and imbalanced data and suffer from inefficiencies due to frequent communication and the aggregation of weights. More critically, the dynamic and complex nature of clinical environments demands scalable AI systems capable of continuously learning from diverse modalities and multilabels. Yet, both centralized and decentralized models are prone to catastrophic forgetting during system expansion, often requiring full model retraining to incorporate new data. To address these limitations, we propose VGS-ATD (Vision Transformer-based Generalized and Scalable Ai-To-Data), a novel distributed learning framework. It leverages fully-connected peer-to-peer communication to ensure privacy without relying on a central server. It conducts a one-time weight aggregation between nodes to maintain domain-specific knowledge, thereby reducing the aggregation and communication burden and preventing catastrophic forgetting. To validate VGS-ATD, we evaluate it on highly heterogeneous and imbalanced medical datasets encompassing a wide range of imaging modalities and classification tasks. In experiments spanning 30 datasets and 80 independent labels across distributed nodes, VGS-ATD achieved an overall accuracy of 92.7%, outperforming centralized learning (84.9%) and swarm learning (72.99%), while federated learning failed under these conditions due to high requirements on computational resources. VGS-ATD also demonstrated strong scalability, with only a 1% drop in accuracy on existing nodes after expansion, compared to a 20% drop in centralized learning, highlighting its resilience to catastrophic forgetting. Additionally, it reduced computational costs by up to 50% relative to both centralized and swarm learning, confirming its superior efficiency and scalability.
Index Terms:
Decentralized learning, Data privacy, Multi-label learning, Vision transformer, Data heterogeneity, Medical imaging.
I Introduction
@@
Nowadays, the development of AI technology has facilitated the emergence of many advanced architectures, further promoting their applications in the real-world environment. In the healthcare domain, AI-supported medical diagnosis has long been a topic of interest and has demonstrated promising performance. However, the statistical heterogeneity and privacy concerns remain crucial challenges in medical imaging learning. The differences in disease symptoms, regional variations, and imaging techniques may lead to heterogeneity issues, which significantly hinder the generalizability and performance of the medical model. Normally, using traditional Centralized Learning (CL) methods to collect and train with a large dataset can improve the model’s generalizability and mitigate the negative influence of data heterogeneity. However, such learning methods have a severe risk of data leakage during model transferring and storage, which has prevented their utilization in many privacy-sensitive areas, especially in the medical field.
@@
To reduce the privacy risk and enable secure sharing between clinics, recent research has introduced two novel decentralized learning methods: Federated Learning (FL) [1] and Swarm Learning (SL) [2, 3], providing a viable option for training models without gathering data centrally. These methods train models locally on client devices and aggregate only the models’ weights to share learned knowledge directly without transferring raw data. Specifically, FL uses a central server to communicate with client nodes, and allows a weight aggregation during the node’s model training. In contrast, the SL method employs a sequential peer-to-peer model update process, passing trained model weights from a previous client to a new client and updating only part of its layers. This method can eliminate the requirement for a central server, reducing communication costs, and allows the model to learn more domain-specific knowledge, thereby reducing the influence of data heterogeneity between clients. However, both FL and SL methods have several limitations, including the inability to handle data with different feature spaces, high communication costs during model aggregation, catastrophic forgetting after updating the global model, and high computational resource requirements for edge device training. Moreover, their solutions aggregate nodes only during the training process, without supporting a free-in and free-out mode for clients who want to enter or exit the system, requiring the entire system to be retrained each time. Moreover, the sequential training process of SL may also compromise the robustness of the entire system, as a single node’s compromise or malicious behavior can affect the integrity of the entire system.
@@
Recently, some novel frameworks have been designed to address the aforementioned challenges, with few providing holistic solutions that simultaneously tackle multiple limitations within a unified architecture. For example, Chakshu’s Orbital Learning framework [4] integrates Split Learning [5] and ensemble learning to achieve both secure and robust performance. It demonstrates improved accuracy over traditional FL methods and supports datasets with varying feature spaces. However, its approach to data heterogeneity is limited, as it merely employs a similarity-based mechanism to avoid aggregating models from nodes with highly divergent features, without truly resolving the underlying heterogeneity issue. Similarly, Thapa et al. proposed SplitFed [6], a hybrid framework that combines the advantages of split learning and FL. This approach mitigates the sequential dependency inherent in traditional split learning and has shown promising performance across several benchmark medical datasets. Nonetheless, their work does not account for complex real-world scenarios such as multi-label classification.
@@
Figure 1: a, Sample of cloud-based centralised learning paradigm, each client’s data and learned weights are stored in the central server. b, Sample of federated learning paradigm, as only learned weights are shared in the central server. Raw data and computational resources are kept by clients. c, Sample of swarm learning process, with both data and weight kept locally with clients. d, Sample of ATD learning process, client can train a local model and share learned weights with each other, and keep the raw data locally.
@@
In this study, we propose a novel learning framework: VGS-ATD, which utilizes the AI-To-Data (ATD) [7] learning method to address several challenges, including data privacy, heterogeneity and imbalance, and computational efficiency, as well as system adaptability and scalability. The ATD is a highly distributed peer-to-peer decentralized learning method, which allows it to exchange the backbone weights of client models and develop a customized global model to handle different scenarios. As shown in Fig. 1, the primary difference between CL, FL, SL, and ATD lies in the method we used, which enables participating nodes to directly connect with other nodes, allowing for flexible weight exchange and dynamic collaboration between clients. Benefiting from ATD’s highly distributed structure, the VGS-ATD conducts a one-time aggregation within each client by exchanging only the model’s body (extractor) weight to build a global extractor first, followed by building a global classification head (classifier) using extracted CLS token embeddings through machine learning classifier training and ensemble learning.
@@
In VGS-ATD, we employ Vision Transformer (ViT) [8] as the foundational model, replacing commonly used Convolutional Neural Networks (CNNs). ViT is a novel deep learning architecture that leverages self-attention mechanisms to capture long-range dependencies and global contextual information. Compared to CNNs, ViT exhibits better performance in complex and high-variance medical imaging tasks, but requires a larger dataset for training to ensure satisfactory performance. Moreover, our method introduces a fully decentralized and distributed learning framework that allows clients to join dynamically without incurring additional overhead. We applied a one-time weight aggregation after local model training finishes, which enables them to learn personalized local feature information to against data heterogeneity. The follow-up feature selection, classifier training, and ensemble learning process leverages multiple robust classifiers, enabling reliable predictions and reducing the influence of feature bias and class imbalance.
@@ II Backgrounds
In the long term, dealing with complex and changeable medical imaging-related tasks remains a challenge to DL models. Multiple issues, including data heterogeneity, data scarcity, privacy concerns, and class imbalance, have hindered the model’s ability to learn powerful feature representations from the target data. So far, both the traditional centralized learning method and the newly emerged decentralized learning methods have their shortcomings and cannot fully handle the complex and computationally constrained medical environment.
II-A Centralized Learning Method
@@
Centralized learning is one of the earliest and most widely adopted paradigms in machine learning, where all data is collected and processed on a central server. Since the success of large-scale CNN architectures such as AlexNet [9] and Xception [10], the CL method has become widely used for training deep learning models. By combining all data and computational resources, the CL method can maximize training effectiveness and data usability, ensuring satisfactory performance, especially with data-hungry architectures. In the medical imaging field, CL methods have shown promising results in maintaining high-level model performance and robustness. However, in a real-world scenario, the availability of data in a single hospital is limited due to the high expense of collecting and annotating it. Uploading medical datasets from hospitals worldwide would require numerous permissions. Moreover, the various requirements from different governments for storing and utilizing trained models make it impossible to create such an environment for conducting CL [11]. Another major challenge is the constraints of computational resources in training with a large amount of medical samples, as large-scale DL architectures can be resource-intensive and time-consuming [12]. Current DL models lack the capability to handle dispersed and heterogeneous medical data, and combining different medical datasets may also lead to differences in feature distribution, causing bias [13]. So far, a widely applied solution is to retrain the model when facing new disease data, which wastes computational resources and is time-consuming. Currently, most clinical data are collected and stored independently within hospitals, and uploading or exchanging raw data poses a high risk of data leakage.
II-B Federated Learning Method
@@
Federated Learning [1] has become a widely adopted paradigm in medical image analysis due to its strong privacy-preserving properties. FL typically falls into three main categories: 1. Horizontal federated learning (HFL): HFL is defined for client datasets that share a homogeneous feature space, but are distinct in the sample ID space. 2. Vertical federated learning (VFL): VFL deals with scenarios where two datasets share the same sample identity space, but differ concerning their feature spaces. 3. Federated transfer learning (FTL): FTL is employed when two datasets vary within both sample ID space and feature space, using transfer learning to connect different clients’ datasets [14]. HFL has gained early popularity for its simplicity and ability to train a global model by aggregating local updates, without sharing raw data. All three approaches offer privacy protection by exchanging only model parameters instead of sensitive data. However, a major limitation of FL lies in its difficulty in handling heterogeneous data distributions, which is a common scenario in real-world applications. This heterogeneity can arise from differences in statistical distributions across participants [15], or from domain shift, where local datasets vary significantly in modality, structure, or class imbalance [16]. Such discrepancies often lead to degraded model performance and hinder generalization across clients.
@@
To address the data heterogeneity problem, several modified FL structures have been proposed, including Federated Representation Learning (FRL) [17] and FedBABU [18], among others. The primary difference between these methods and general FL methods lies in their decomposition of the entire network into two components: the body (extractor), which is related to universality, and the head (classifier), which is related to personalization. The weight aggregation would only happen at the extractor level, and each client node trains a personalized classifier after receiving the aggregated global extractor from the central server. In this way, FRL and FedBABU achieved success in allowing the decentralized FL paradigm to learn from clients with heterogeneous feature spaces and distinct sample ID spaces. Nevertheless, such methods require additional training resources on the client side and do not allow any new client node to participate or an existing client node to quit the system, as it necessitates the cooperation of all remaining clients to produce a new aggregated extractor.
II-C Swarm Learning Method
@@
Swarm learning [2, 3] emerges as a decentralized paradigm that promotes collaborative learning through peer-to-peer interaction, utilizing edge computing and blockchain technology. In contrast to many existing federated learning models, SL’s P2P structure surpasses them, eliminating the need for a central server. Moreover, the blockchain technology offers strong guarantees regarding data privacy and secure communication among the nodes [19]. Warnat’s work [2], which utilizes SL in medical imaging tasks, achieved comparable performance in chest X-ray and blood cell classification tasks compared to centralized learning techniques. So far, SL has demonstrated its potential in addressing the limitations of centralized data storage and model aggregation. While SL provides a robust defense against adversarial attacks, it comes at a high computational expense. Another main limitation of SL is that its sequential peer-to-peer structure is inherently time-consuming, which does not support parallel learning like FL. Even worse, the whole system may fail if a single node becomes compromised or acts maliciously.
II-D AI To Data Method
@@
AI-To-Data [7] is a newly proposed decentralized learning method developed to address the challenges of previous decentralized learning methods like FL and SL. The client nodes within the ATD system function as both local and global nodes, and they can exchange trained weights directly between each other without going through a central server. ATD applied a Peer-to-Peer (P2P) blockchain technique to ensure security during node interaction and communication. Benefit from P2P communication, as each node can aggregate weights locally and build a customized global model tailored to its needs, even if another node in the system encounters issues. Moreover, ATD supports a flexible training option with both parallel and sequential approaches, as the weight aggregation in ATD occurs after each node trains its local model. Under this structure, each client node can train a local model without interrupting other nodes and also utilize the weights shared from other nodes to perform transfer learning, thereby boosting training efficiency. For models trained on different feature spaces and sample identity spaces, ATD enables them to aggregate only the backbone extractors and train a global classifier for prediction. Compared to the FL method, ATD eliminated the need for a central server, reducing the cost of node-central server communication and the risk of privacy leakage during communication. In contrast to SL, P2P communication between nodes facilitates model training in a more resilient and flexible manner, ensuring the system’s stability and avoiding the chain reaction in SL, where one node’s failure leads to the entire system failing. So far, the effectiveness of ATD has only been proven in surveillance videos [20], using CNNs as the base models; more domains and architectures need to be tested to further boost its development.
@@
TABLE I: An abstract comparison of Centralized learning (CL), Federated learning (FL), Swarm learning (SL), and AI to Data (ATD).
@@
In summary, as AI systems are increasingly involved in medical applications, the standards and requirements for next-generation medical AI systems are becoming clearer. The need for a more powerful, generalized, scalable, stable, and privacy-preserved learning method is growing. However, as Table I shows, existing methods like CL, FL, and SL have limitations in handling large-scale multi-label medical classification tasks. A more distributed AI-to-data method, on the other hand, has demonstrated promising potential in addressing these challenges.
III Methodology

### v3

@@ Abstract
In recent years, advanced deep learning architectures have shown strong performance in medical imaging tasks. However, the traditional centralized learning paradigm poses serious privacy risks as all data is collected and trained on a single server. To mitigate this challenge, decentralized approaches such as federated learning and swarm learning have emerged, allowing model training on local nodes while sharing only model weights. While these methods enhance privacy, they struggle with heterogeneous and imbalanced data and suffer from inefficiencies due to frequent communication and the aggregation of weights. More critically, the dynamic and complex nature of clinical environments demands scalable AI systems capable of continuously learning from diverse modalities and multilabels. Yet, both centralized and decentralized models are prone to catastrophic forgetting during system expansion, often requiring full model retraining to incorporate new data. To address these limitations, we propose VGS-ATD (Vision Transformer-based Generalized and Scalable AI-To-Data), a novel distributed learning framework. It leverages fully-connected peer-to-peer (P2P) communication to ensure privacy without relying on a central server. It conducts a one-time weight aggregation between nodes to maintain domain-specific knowledge, thereby reducing the aggregation and communication burden and preventing catastrophic forgetting. To validate VGS-ATD, we evaluate it on highly heterogeneous and imbalanced medical datasets encompassing a wide range of imaging modalities and classification tasks. In experiments spanning 30 datasets and 80 independent labels across distributed nodes, VGS-ATD achieved an overall accuracy of 92.7%, outperforming centralized learning (84.9%) and swarm learning (72.9%), while federated learning failed under these conditions due to high requirements on computational resources. VGS-ATD also demonstrated strong scalability, with only 1% drop in accuracy on existing nodes after expansion, compared to a 20% drop in centralized learning, highlighting its resilience to catastrophic forgetting. Additionally, it reduced computational costs by up to 50% relative to both centralized and swarm learning, confirming its superior efficiency and scalability.
keywords:
[…]
1 Introduction
@@
Nowadays, the development of AI technology has facilitated the emergence of numerous advanced architectures, further promoting their applications in real-world environments [beddiar2023deep]. In the healthcare domain, AI-supported medical diagnosis has long been a topic of interest and has demonstrated promising performance [beddiar2023deep, liu2025adaptive]. However, the statistical heterogeneity and privacy concerns remain crucial challenges in medical imaging learning. The differences in disease symptoms, regional variations, and imaging techniques may lead to heterogeneity issues, which significantly hinder the generalizability and performance of the medical model. Normally, using traditional Centralized Learning (CL) methods to collect and train with a large dataset can improve the model’s generalizability and mitigate the negative influence of data heterogeneity. However, such learning methods have a severe risk of data leakage during model transferring and storage, which has prevented their utilization in many privacy-sensitive areas, especially in the medical field [beltran2023decentralized].
@@
To reduce the privacy risk and enable secure sharing between clinics, recent research has introduced two novel decentralized learning methods: Federated Learning (FL) [mcmahan2017communication] and Swarm Learning (SL) [warnat2021swarm, saldanha2022swarm], providing a viable option for training models without gathering data centrally. These methods train models locally on client devices and aggregate only the models’ weights to share learned knowledge directly without transferring raw data. Specifically, FL uses a central server to communicate with client nodes and conduct weight aggregation to achieve a more powerful and generalized global model [banabilah2022federated]. In contrast, the SL method employs a sequential P2P model update process, passing trained model weights from a previous client to a new client and conducting model aggregation with the node’s local model [shammar2024swarm]. However, both FL and SL methods have several limitations, including the inability to handle data heterogeneity, high communication costs during model aggregation, catastrophic forgetting after updating the global model, and high computational resource requirements for training on edge devices [rauniyar2023federated]. Their solutions aggregate a global model during the node local model training process, without supporting a free-in and free-out mode for clients who want to enter or exit the system. Moreover, the sequential training process of SL may also compromise the robustness of the entire system, as a single node’s compromise or malicious behavior can affect the integrity of the entire system.
@@
Recently, some novel frameworks have been designed to address the aforementioned challenges, providing holistic solutions that simultaneously tackle multiple limitations within a unified architecture. For example, Chakshu’s Orbital Learning framework [chakshu2024orbital] integrates Split Learning [vepakomma2018split] and ensemble learning to protect data privacy and enhance model performance. It demonstrates improved accuracy over traditional FL methods and supports datasets with varying feature spaces. However, its approach to data heterogeneity is limited, as it merely employs a similarity-based mechanism to avoid aggregating models from nodes with highly divergent features. Similarly, Thapa et al. proposed SplitFed [thapa2022splitfed], a hybrid framework that combines the advantages of split learning and FL. This approach mitigates the sequential dependency inherent in traditional split learning and has shown promising performance across several benchmark medical datasets. Nonetheless, their work does not account for complex real-world scenarios such as multi-label classification.
@@
Figure 1: a, Sample of cloud-based centralized learning paradigm, each client’s data and learned weights are stored in the central server. b, Sample of federated learning paradigm, as only learned weights are shared in the central server. Raw data and computational resources are kept by clients. c, Sample of swarm learning process, with both data and weight kept locally with clients. d, Sample of ATD learning process, the client can train a local model and share learned weights with each other, and keep the raw data locally. Unlike others, ATD operates efficiently at both intra-node and inter-node levels: within each node, data is split into low-resource batches (trained on single or multi-GPU setups) and trained into a local model. These local models are then shared in a fully connected manner, enabling collaboration without central coordination. ATD supports continuous, incremental learning as new data or tasks emerge, while preserving privacy and minimizing costly retraining.
@@
In this study, we propose a novel learning framework: Vision Transformer-based Generalized and Scalable AI-To-Data (VGS-ATD), which utilizes the AI-To-Data (ATD) [alzubaidi2025atd] learning method to address several challenges, including data privacy, data heterogeneity and imbalance, and computational efficiency, as well as system adaptability and scalability. The ATD is a highly distributed P2P decentralized learning method, which allows nodes to exchange local model weights and develop a customized global model to handle different scenarios. As shown in Fig. 1, the primary difference between CL, FL, SL, and ATD lies in the system structure used, as ATD enables participating nodes to directly connect with other nodes, allowing for more flexible weight exchange and dynamic collaboration between clients. Benefiting from ATD’s highly distributed structure, the proposed VGS-ATD framework conducts a one-time aggregation within each client by exchanging only the model’s body (extractor) weight to build a global extractor first, followed by building a global classification head (classifier) using extracted CLS token embeddings through machine learning classifier training and ensemble learning.
@@
In VGS-ATD, we employ Vision Transformer (ViT) [dosovitskiy2020image] as the foundational model, replacing commonly used Convolutional Neural Networks (CNNs). ViT is a novel deep learning architecture that leverages self-attention mechanisms to capture long-range dependencies and global contextual information. Compared to CNNs, ViT exhibits better performance in complex and high-variance medical imaging tasks, but requires a larger dataset for training to ensure satisfactory performance. Moreover, our method introduces a fully decentralized and distributed learning framework that allows clients to join dynamically without incurring additional overhead. We applied a one-time weight aggregation after local model training finishes, which enables them to learn personalized local feature information to against data heterogeneity. The follow-up feature selection, classifier training, and ensemble learning process leverages multiple robust classifiers, enabling reliable predictions and reducing the influence of feature bias and class imbalance.
@@ 2 Backgrounds
Dealing with complex and changeable medical imaging-related tasks remains a challenge to DL models. Multiple issues, including data heterogeneity, data scarcity, privacy concerns, and class imbalance, have hindered the model’s ability to learn powerful feature representations from the target data. In this section, we introduce several centralized and decentralized methods that appear in the proposed method or as baselines, along with the one-time aggregation technique we apply for instant model aggregation.
2.1 Centralized Learning Method
@@
Centralized learning is one of the earliest and most widely adopted paradigms in machine learning, where all data is collected and processed on a central server. Since the success of large-scale CNN architectures such as AlexNet [krizhevsky2017imagenet] and Xception [chollet2017xception], the CL method has become widely used for training deep learning models. By combining all data and computational resources, the CL method can maximize training effectiveness and data usability, ensuring satisfactory performance, especially with data-hungry architectures. In the medical imaging field, CL methods have shown promising results in maintaining high-level model performance and robustness. However, in a real-world scenario, the availability of data in a single hospital is limited due to the high expense of collecting and annotating it. Uploading medical datasets from hospitals worldwide would require numerous permissions. Moreover, the various requirements from different governments for storing and utilizing trained models make it impossible to create such an environment for conducting CL [bao2022federated]. Another major challenge is the constraints of computational resources in training with a large amount of medical samples, as large-scale DL architectures can be resource-intensive and time-consuming [bellavista2021decentralised]. Current DL models lack the capability to handle dispersed and heterogeneous medical data, and combining different medical datasets may also lead to differences in feature distribution, causing bias [yue2020deep]. So far, a widely applied solution is to finetune part of the model when facing new disease data, which wastes computational resources and is time-consuming. Additionally, most clinical data are collected and stored independently within hospital servers, and uploading or exchanging raw data poses a high risk of data leakage.
2.2 Federated Learning Method
@@
Federated Learning [mcmahan2017communication, chen2025prototype, sultan2025federated] has become a widely adopted paradigm in medical image analysis due to its strong privacy-preserving properties. FL typically falls into three main categories: 1. Horizontal federated learning (HFL): HFL is defined for client datasets that share a homogeneous feature space, but are distinct in the sample ID space. 2. Vertical federated learning (VFL): VFL deals with scenarios where two datasets share the same sample identity space, but differ concerning their feature spaces. 3. Federated transfer learning (FTL): FTL is employed when two datasets vary within both sample ID space and feature space, using transfer learning to connect different clients’ datasets [guendouzi2023systematic]. HFL has gained early popularity for its simplicity and ability to train a global model by aggregating local updates, without sharing raw data. All three approaches offer privacy protection by exchanging only model parameters instead of sensitive data. However, a major limitation of FL lies in its difficulty in handling heterogeneous data distributions, which is a common scenario in real-world applications. This heterogeneity can arise from differences in statistical distributions across participants [gao2022survey], or from domain shift, where local datasets vary significantly in modality, structure, or class imbalance [stacke2020measuring]. Such discrepancies often lead to degraded model performance and hinder generalization across clients.
@@
To address the data heterogeneity problem, several modified FL structures have been proposed, including Federated Representation Learning (FRL) [jing2023exploring] and FedBABU [oh2021fedbabu]. The primary difference between these methods and general FL methods lies in their decomposition of the entire network into two components: the body (extractor), which is related to universality, and the head (classifier), which is related to personalization. The weight aggregation would only happen at the extractor level, and each client node trains a personalized classifier after receiving the aggregated global extractor from the central server. In this way, FRL and FedBABU achieved success in allowing the decentralized FL paradigm to learn from clients with heterogeneous feature spaces and distinct sample ID spaces. Nevertheless, such methods require additional training resources on the client side and do not allow any new client node to instantly participate or quit the system, as it necessitates the cooperation of all remaining clients to produce a new aggregated extractor.
2.3 Swarm Learning Method
@@
Swarm learning [warnat2021swarm, saldanha2022swarm] emerges as a decentralized paradigm that promotes collaborative learning through P2P interaction, utilizing edge computing and blockchain technology. In contrast to many existing federated learning models, SL’s P2P structure surpasses them, eliminating the need for a central server. Moreover, the blockchain technology offers strong guarantees regarding data privacy and secure communication among the nodes [fragkou2024joint]. Warnat’s work [warnat2021swarm], which utilizes SL in medical imaging tasks, achieved comparable performance in chest X-ray and blood cell classification tasks compared to centralized learning techniques. So far, SL has demonstrated its potential in addressing the limitations of centralized data storage and model aggregation. While SL provides a robust defense against adversarial attacks, it comes at a high computational expense. Another main limitation of SL is that its sequential P2P structure is inherently time-consuming, which does not support parallel learning like FL. Even worse, the whole system may fail in a chain reaction if a single node becomes compromised or acts maliciously.
2.4 AI To Data Method
@@
AI-To-Data [alzubaidi2025atd] is a newly proposed decentralized learning method developed to address the common challenges mentioned in FL and SL. As AI systems are increasingly involved in medical applications, the need for a more powerful, generalized, scalable, stable, and privacy-preserved next-generation medical AI systems becomes urgent. As Table 1 shows, existing methods like CL, FL, and SL have limitations in handling large-scale multi-label medical classification tasks. On the other hand, the ATD method has demonstrated promising potential in addressing these challenges due to its flexible and modular design. The client nodes within the ATD system function as both local and global nodes, and can exchange trained weights directly without going through a central server. ATD applied a P2P blockchain technique to ensure security during node interaction and communication. Benefit from P2P communication, each node can aggregate weights locally and build a customized global model to follow its personalized needs. Moreover, ATD supports a flexible training option with both parallel and sequential approaches, as the weight aggregation in ATD occurs after the local node model training is complete. Under this structure, each client node can train a local model without interrupting other nodes and also utilize the weights shared from other nodes to perform transfer learning, thereby achieving higher performance and boosting training efficiency. For models trained on different feature spaces and sample identity spaces, ATD enables them to aggregate only the backbone extractors and train a global classifier for prediction. Compared to the FL method, ATD eliminated the need for a central server, reducing the cost of node-central server communication and the risk of privacy leakage during communication. In contrast to SL, a more flexible P2P communication between nodes facilitates model training in a more resilient and flexible manner, ensuring the system’s stability and avoiding the chain failure reaction in SL. So far, the effectiveness of ATD has only been proven in surveillance videos [jebur2025scalable], using CNNs as the base models; more domains and architectures need to be tested to further boost its development.
@@
Table 1: An abstract comparison of Centralized learning (CL), Federated learning (FL), Swarm learning (SL), and AI to Data (ATD).
@@
2.5 One-time Weight Aggregation
One-time weight aggregation is an emerging and promising research direction in decentralized learning that significantly reduces communication costs. In recent FL studies, this approach attracts much attention and is often referred to as one-shot federated learning. For example, Liu et al. proposed FedLPA [liu2024fedlpa], where clients train models locally and compute the local covariance of their datasets using a layer-wise Laplace approximation. The server then aggregates these contributions to estimate the global expectation and update the global model parameters in a single aggregation step. Similarly, Zeng et al. introduced IntactOFL [zeng2024one], which integrates well-trained local models into an ensemble model and uses it as a teacher to guide a student global model. The ensemble, implemented as a mixture-of-experts network, is strengthened with self-supervised generated data and frozen to preserve local knowledge and prevent catastrophic forgetting. This ensemble then guides the training of a high-performance global gating network. Experimental results from both works show that one-time aggregation not only reduces communication overhead but can also outperform methods relying on multiple aggregation rounds, by better capturing global parameters while preserving local knowledge. In this study, we adopt a similar idea by training local models independently and then aggregating them once through P2P node communication, where all parameters are used to maintain local knowledge.
3 Methodology